# Dynamic Programming: From Local Optimality to Global Optimality

### v2

@@ 3.2. Strong Irreducibility
Recall that a linear subspace $I$ of $b\mathsf{X}$ is called an ideal in $b\mathsf{X}$ when $f\in I$ and $|g|\leqslant|f|$ implies $g\in I$. An ideal $I$ is said to be invariant for a linear operator $K$ if $KI\subset I$. A linear operator $K$ from $b\mathsf{X}$ to itself is called positive when $Kf\geqslant 0$ for all $f\geqslant 0$. A positive linear operator $K$ is called irreducible if the only invariant ideals under $K$ are the trivial subspace $\{0\}$ and the whole space $b\mathsf{X}$. We will make use of the characterization in Proposition 8.3 (c) of Schaefer (1974): A positive linear operator $K$ on $b\mathsf{X}$ is irreducible if and only if, for each nonzero $f\in b\mathsf{X}_{+}$ and each nonzero $\mu\in b\mathsf{X}^{\prime}_{+}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mu,K^{m}f}\right\rangle>0$.
In what follows, we call a transition kernel $P$ on $\mathsf{X}$ strongly irreducible if its Markov operator (see (2)) is irreducible on $b\mathsf{X}$ in the sense just defined.
Here is our main result for the strongly irreducible case. In the statement $\sigma$ is any feasible policy.
@@
In view of Lemma 3.1, it suffices to show that (E1) and (E3) are equivalent. That (E3) implies (E1) is immediate from the definition of optimal policies. Hence we need only show that (E1) implies (E3). In line with the conditions of Theorem 3.3, we assume that $\sigma$ is a feasible policy and $P_{\sigma}$ is strongly irreducible.
Let $h\coloneq v^{*}-v_{\sigma}$. By the definition of $v^{*}$ we have $0\leqslant h$. We claim in addition that $h=0$. To see this, suppose to the contrary that $h$ is nonzero. In this case, by strong irreducibility, for each nonzero $\mu$ in the positive cone of $b\mathsf{X}^{\prime}$ we can find an $m\in\mathbbm{N}$ such that $\left\langle{\mu,P_{\sigma}^{m}\,h}\right\rangle>0$. Because $\delta_{\bar{x}}$ is a nonzero element of the positive cone of $b\mathsf{X}^{\prime}$, we can set $\mu=\delta_{\bar{x}}$ to obtain an $m\in\mathbbm{N}$ with $(P_{\sigma}^{m}\,h)(\bar{x})>0$. This contradicts (6), so $h=0$ holds. In other words, $v_{\sigma}(x)=v^{*}(x)$ for all $x\in\mathsf{X}$, as was to be shown. ∎
3.3. Finite States and Actions
In this section, we specialize to the case where $\mathsf{X}$ is finite and study the role of irreducibility in this setting. Note that the previous results (in particular, Theorem 3.3), can be applied by taking the metric on $\mathsf{X}$ to be the discrete metric, so that all subsets of $\mathsf{X}$ are open and $\mathscr{B}$ is the set of all subsets of $\mathsf{X}$. We also assume that $\mathsf{A}$ is finite and impose the discrete metric on $\mathsf{A}$.
Given a transition kernel $Q$ on $\mathsf{X}$, we say that $y$ in $\mathsf{X}$ is $Q$-accessible from $x\in\mathsf{X}$ when there exists an $m\in\mathbbm{N}$ such that $Q^{m}(x,y)>0$. The usual definition of irreducibility of a transition kernel $Q$ on finite $\mathsf{X}$ is that, for every $x,y\in\mathsf{X}$, $x$ is $Q$-accessible from $y$ and $y$ is $Q$-accessible from $x$. To distinguish between different notions of irreducibility, we call this discrete irreducibility.
Lemma 3.4.
  1. (a)
    $P_{\sigma}$ is strongly irreducible.
  2. (b)
    $P_{\sigma}$ is discretely irreducible.
Proof.
((a) $\implies$ (b)) Fix $x,y\in\mathsf{X}$. Let $\mathbbm{1}_{x}(z)$ equal $1$ when $z=0$ and zero elsewhere. Let $\mathbbm{1}_{y}$ be defined analogously. Since $\mathbbm{1}_{x}\in bm\mathsf{X}_{+}$ and $\mathbbm{1}_{y}\in(bm\mathsf{X}_{+})^{\prime}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle>0$. But $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle=P_{\sigma}^{m}(x,y)$, so $y$ is $P_{\sigma}$-accessible from $x$. This proves that $P_{\sigma}$ is discretely irreducible.
((b) $\implies$ (a)) Fix a nonzero $f\in bm\mathsf{X}_{+}$ and nonzero $\mu\in(bm\mathsf{X}_{+})^{\prime}$. Since $\mathsf{X}$ is finite, these element are just maps from $\mathsf{X}$ to $\mathbbm{R}_{+}$ and, as both are nonzero, we can find $\bar{x},\bar{y}\in\mathsf{X}$ such that $f(\bar{y})>0$ and $\mu(\bar{x})>0$. Moreover, since $P_{\sigma}$ is discretely irreducible, there is $m\in\mathbbm{N}$, such that $P_{\sigma}^{m}(\bar{x},\bar{y})>0$. As a result,
This proves that $P_{\sigma}$ is strongly irreducible. ∎
We can now state a result for the discrete case, where $\sigma$ is any feasible policy. The result is immediate from Lemma 3.4 and Theorem 3.3.
Corollary 3.5.
If $P_{\sigma}$ is discretely irreducible, then (E1)(E3) are equivalent.
3.4. The Significance of Irreducibility
In this section, we show that the irreducibility assumptions used in Theorems 3.3 cannot be dropped: without irreducibility, (E1)(E3) are not generally equivalent. To show this, we consider a two-state MDP with $\mathsf{X}=\{1,2\}$ and $\mathsf{A}=\{1,2\}$. The feasible coprrespondence is defined by $\Gamma(1)=\{1,2\}$ and $\Gamma(2)=\{2\}$. The reward function is defined by $r(i,j)=r_{ij}$ with
$P_{\sigma}=P_{\pi}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
Now we compute the lifetime value functions for the optimal $\sigma$ and $\pi$. For policy $\sigma$, we have
$r_{\sigma}=(r_{12},r_{22})=(1,2),\quad v_{\sigma}=(I-\beta P_{\sigma})^{-1}r_{\sigma}=(10,20).$ (9)
For $\pi$, we have
$r_{\pi}=(r_{11},r_{22})=(0,2),\quad v_{\pi}=(I-\beta P_{\pi})^{-1}r_{\pi}=(0,20).$ (10)
Since, $v^{*}(x)\coloneq\sup_{s\in\Sigma}v_{s}(x)=v_{\sigma}(x)$, $\sigma$ is an optimal policy. On the other hand, $v_{\pi}(2)=v^{*}(2)$, but $v_{\pi}(1)<v^{*}(1)$, indicating that optimality in one state does not guarantee global optimality. This shows that irreducibility cannot be dropped from the statement of Theorem 3.3.
@@ 4.1. Reachable States
As a preliminary step, we show that $h\coloneq v^{*}-v_{\sigma}$ is continuous under the stated assumptions. Since $\sigma$ is continuous, our conditions on $(r,\Gamma,\beta,P)$ imply that the mappings $x\mapsto\int v(x^{\prime})P(x,\sigma(x),dx^{\prime})$ and $x\mapsto r(x,\sigma(x))$ are continuous on $\mathsf{X}$ whenever $v\in bc\mathsf{X}$. This implies that $T_{\sigma}$ is invariant on $bc\mathsf{X}$. Moreover, $bc\mathsf{X}$ is a closed subset of the complete metric space $bm\mathsf{X}$ under the supremum norm (since uniform limits of continuous functions are continuous). In addition, given $v,w\in bm\mathsf{X}$, we have
@@
Since $\beta<1$, the contraction mapping theorem implies that $T_{\sigma}^{n}w\to v_{\sigma}$ for every $w\in bm\mathsf{X}$. If we now fix $w\in bc\mathsf{X}$ and use the fact that $T_{\sigma}$ is invariant on this set, we obtain a sequence $(T^{n}w)_{n\in\mathbbm{N}}$ converging to $v_{\sigma}$ and entirely contained in $bc\mathsf{X}$. As $bc\mathsf{X}$ is closed in $bm\mathsf{X}$, this implies that $v_{\sigma}$ is in $bc\mathsf{X}$. In particular, $v_{\sigma}$ is continuous. As $v^{*}$ is also continuous (see Proposition 2.2), we see that $h$ is continuous.
@@
But $v_{\sigma}(x)=v^{*}(x)$, so this inequality contradicts Lemma 2.3. The contradiction proves Theorem 4.1. ∎
Let us briefly consider how this translates to MDPs with finite state and action spaces. We give $\mathsf{X}$ and $\mathsf{A}$ the discrete topology, under which every set is open. In this setting, $y\in$ is $Q$-reachable from $x\in\mathsf{X}$ if and only if $y$ is $Q$-accessible from $x$. Moreover, every function from $\mathsf{X}$ to $\mathsf{A}$ is continuous. These observations lead to the next corollary.
Corollary 4.2.
If $v_{\sigma}(x)=v^{*}(x)$ and $y$ is $P_{\sigma}$-accessible from $x$, then $v_{\sigma}(y)=v^{*}(y)$
@@ 4.2. Open Set Irreducibility
In view of Lemma 3.1, it suffices to show (E1) implies (E3). So fix $x\in\mathsf{X}$ and suppose that $v_{\sigma}(x)=v^{*}(x)$. For any $y\in\mathsf{X}$, open set irreducibility implies that $y$ is $P_{\sigma}$-reachable from $x$. Hence, by Theorem 4.1, we have $v_{\sigma}(y)=v^{*}(y)$. In particular, (E3) holds. ∎
@@ 4.3. $\pi$-Irreducibility
Now we state a result for the weakly irreducible case. In the statement $\sigma$ is any feasible policy.
@@
5. Application: Optimal Saving with Stochastic Returns
In this section, we examine an optimal savings problem with stochastic returns on wealth. Our aim is to illustrate the theoretical results stated above. In the problem, an agent seeks to maximize lifetime utility by choosing an optimal consumption plan. The evolution of wealth is governed by the equation
@@ 5. Application: Optimal Saving with Stochastic Returns
We formulate this problem as an MDP. The state space is $\mathbbm{R}_{+}$ and the set of feasible actions at wealth level $w$ is $\Gamma(w)=\{c\in\mathbbm{R}_{+}:c\leqslant w\}$. A feasible policy in this setup is a Borel measurable function $\sigma$ from $\mathbbm{R}_{+}$ to itself satisfying $\sigma(w)\leqslant w$ for all $w\in\mathbbm{R}_{+}$. The reward function is $r(w,c)\coloneq u(c)$, where $u(c)$ is the utility derived from consumption and $u$ is continuous and strictly concave on $\mathbbm{R}_{+}$.
@@
Let $\Sigma$ be the set of feasible policies. Fix $w\in\mathbbm{R}_{+}$ and an open set $B\subseteq\mathbbm{R}_{+}$ with $\pi(B)>0$. Let $\alpha=w-\sigma(w)$. By a change of variable, we obtain
@@ 5.1. Computation
Given the open set irreducibility of the transition kernel at any feasible policy stated in Lemma 5.1, Theorem 4.3 implies that we can compute a (globally) optimal policy by maximizing $v_{\sigma}(w)$ at any fixed $w\in\mathbbm{R}_{+}$. We now explore this result in a computational experiment, where the maximization of $v_{\sigma}(w)$ is based on the Deep Deterministic Policy Gradient (DDPG) algorithm of Lillicrap et al. (2019). The DDPG algorithm uses an actor-critic (AC) method approach that iteratively improves both the policy guess and the value guess (Witten, 1977; Barto et al., 1983; Mnih et al., 2016; Wu et al., 2017; Sutton and Barto, 2018).
Our implementation uses neural networks to approximate both the optimal policy $\sigma$ and the value function $v$. Let $\theta,\lambda\in\mathbbm{R}^{d}$ be the network parameters that characterize the policy and value networks respectively. The policy network $\hat{\sigma}(\cdot;\theta)$ approximates $\sigma$, while the value network $\hat{v}_{\sigma}(\cdot;\lambda)$ approximates $v_{\sigma}$.
Let $N$ be the batch size of Monte Carlo samples. In each training episode, we draw two sequences of wealth levels: current wealth $(w_{i})_{i=1}^{N}$ and corresponding next-period wealth $(w_{i}^{\prime})_{i=1}^{N}$, where successive pairs $(w_{i},w_{i}^{\prime})$ evolve according to the wealth dynamics (11). The parameter updates in each episode of the training run as follows:
  1. (a)
    Given value parameter vector $\lambda$, the policy improvement updates the policy network parameters $\theta$ to maximize the expected return
@@
    $J(\theta)=\frac{1}{N}\sum_{i=1}^{N}u(\hat{\sigma}(w_{i};\theta))+\beta\hat{v}_{\sigma}(w^{\prime}_{i};\lambda).$ (13)
  2. (b)
    Given policy network parameter vector $\theta$, the policy evaluation updates the value network parameters $\lambda$ to minimize the loss function
    $L(\lambda)=\frac{1}{N}\sum_{i=1}^{N}\bigg{[}\hat{v}_{\sigma}(w_{i};\lambda)-\underbrace{(u(\hat{\sigma}(w_{i};\theta))+\beta\hat{v}_{\sigma}(w^{\prime}_{i};\lambda))}_{\text{target value}}\bigg{]}^{2}.$ (14)
Full details of the algorithm are given in Algorithm 1 in Appendix. The policy network can be viewed as an actor who estimates the policy to maximize expected lifetime return, while the value network is a critic who evaluates the policy and guides the actor by providing the continuation value. The algorithm iteratively updates $\theta$ and $\lambda$ to find the optimal policy guided by a constantly improving value function.
At each iteration, the policy network aims to obtain a policy that maximizes an estimate of the $\sigma$-value function as provided by
$\hat{v}_{\sigma}(w;\lambda)=r_{\hat{\sigma}}(w)+\beta\mathbbm{E}_{w^{\prime}}[\hat{v}_{\sigma}(w^{\prime};\lambda)],$ (15)
where $r_{\hat{\sigma}}(w)=u(\hat{\sigma}(w;\theta))$. Through recursive substitution of the approximated $\sigma-$value function, we obtain
$\hat{v}_{\sigma}(w;\lambda)$ $=r_{\hat{\sigma}}(w)+\beta\mathbbm{E}_{w^{\prime}}[r_{\hat{\sigma}}(w^{\prime})+\beta\mathbbm{E}_{w^{\prime\prime}}[\hat{v}_{\sigma}(w^{\prime\prime};\lambda)]]$
$=r_{\hat{\sigma}}(w)+\beta P_{\hat{\sigma}}r_{\hat{\sigma}}(w)+\beta^{2}(P_{\hat{\sigma}})^{2}\hat{v}_{\sigma}(w;\lambda).$ (16)
Continuing this substitution yields
$\hat{v}_{\sigma}(w;\lambda)=\sum_{t=0}^{\infty}\beta^{t}(P_{\hat{\sigma}})^{t}r_{\hat{\sigma}}(w),$ (17)
which corresponds to the lifetime value of policy $\sigma$ in equation (3) evaluated at state $w\in\mathsf{W}$. Thus, our algorithm provides a practical implementation of policy optimization at a single state $w\in\mathsf{W}$.
By Lemma 5.1, the transition kernel $P_{\sigma}$ in our optimal saving problem exhibits open set irreducibility for all feasible policies $\sigma\in\Sigma$. Moreover, Proposition 5.2 ensures that the optimal policy is unique and continuous. Together with Theorem 4.3, the optimal policy can be derived from any wealth level $w\in\mathbbm{R}_{+}$ using Algorithm 1.
Let the value function obtained by following the policy network $\hat{\sigma}$ be $v_{\hat{\sigma}}$.
To verify our theory, we benchmark the result from optimizing the policy at one state against Optimistic Policy Iteration (OPI), a variant of Value Function Iteration (VFI) that is known to converge globally to the optimal policy in this model-based setting (Sargent and Stachurski, 2025). The OPI algorithm operates on a discretized state space and serves as our ground truth comparison. Specifically, we compute $v^{*}$ by applying OPI over a fine grid of wealth levels, which yields a global approximation of the optimal value function and corresponding optimal policy $\sigma^{*}$.
In contrast, our method computes $v_{\hat{\sigma}}$ by fixing a single initial wealth level $w_{0}\in\mathsf{W}$ and maximizing the value network ouput $\hat{v}_{\sigma}(w_{0})$ over the set of policies $\sigma\in\Sigma$, parameterized by the policy network vector $\theta$. We then take the resulting policy $\hat{\sigma}$ and, holding this policy fixed, calculate the entire function $v_{\hat{\sigma}}$ (or, more correctly, the value of the function at different wealth levels on a grid) by computing the lifetime value of $\hat{\sigma}$ from alternative initial conditions. Finally, we compare $v_{\hat{\sigma}}$ with the globally optimal solution $v^{*}$.
Figure 1. $v_{\hat{\sigma}}$ and $\hat{\sigma}$ with $w=50$ against the OPI solutions.
@@
Figure 2. Expected returns (13) and value loss (14) over training episode for the irreducible optimal saving model at $w=50$.
@@
Figure 1 shows the result of these computations when $w_{0}=50$. The function $v_{\hat{\sigma}}$, shown in blue, closely matches the globally optimal value function $v^{*}$ computed via OPI (red dotted line). Similarly, in the second panel, the approximated policy $\hat{\sigma}$ (blue line) closely matches the optimal policy $\sigma^{*}$ (red dotted line). This convergence demonstrates that both the $\hat{\sigma}$-value function $v_{\hat{\sigma}}$ and the policy $\hat{\sigma}$ successfully recover their globally optimal counterparts $v^{*}$ and $\sigma^{*}$, respectively.
@@
In subsequent experiments, we tested the robustness of this outcome to variation in the fixed initial condition $w_{0}$. We found that, as predicted by theory, the resulting function $v_{\hat{\sigma}}$ again closely approximates $v^{*}$, and the resulting policy $\hat{\sigma}$ again closely approximates $\sigma^{*}$. See Figure 6 in Appendix B for a visualization of one experiment.
The training dynamics, illustrated in Figure 2, show consistent improvement in both performance metrics. The expected returns (13) exhibit steady increase, while the value loss (14) decreases throughout the training episodes. The training process stabilizes over episodes. Moreover, the value loss consistently remains below the levels observed in the reducible cases presented in Figures 5(a) and 5(b) of the following section.
5.2. Reducible optimal savings MDP

### v3

@@ 3.2. Strong Irreducibility
Recall that a linear subspace $I$ of $b\mathsf{X}$ is called an ideal in $b\mathsf{X}$ when $f\in I$ and $|g|\leqslant|f|$ implies $g\in I$. An ideal $I$ is said to be invariant for a linear operator $K$ if $KI\subset I$. A linear operator $K$ from $b\mathsf{X}$ to itself is called positive when $Kf\geqslant 0$ for all $f\geqslant 0$. A positive linear operator $K$ is called irreducible if the only invariant ideals under $K$ are the trivial subspace $\{0\}$ and the whole space $b\mathsf{X}$. By Proposition 8.3 (c) of Schaefer (1974), such a $K$ is irreducible if and only if, for each nonzero $f\in b\mathsf{X}_{+}$ and each nonzero $\mu\in b\mathsf{X}^{\prime}_{+}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mu,K^{m}f}\right\rangle>0$.
In what follows, we call a transition kernel $P$ on $\mathsf{X}$ strongly irreducible if its Markov operator (see (2)) is irreducible on $b\mathsf{X}$ in the sense just defined. Here is our main result for the strongly irreducible case. In the statement, $\sigma$ is any feasible policy.
@@
In view of Lemma 3.1, it suffices to show that (E1) and (E3) are equivalent. That (E3) implies (E1) is immediate from the definition of optimal policies. Hence we need only show that (E1) implies (E3). In line with the conditions of Theorem 3.3, we assume that $\sigma$ is a feasible policy and $P_{\sigma}$ is strongly irreducible. Using (E1), we take $x\in\mathsf{X}$ with $v_{\sigma}(x)=v^{*}(x)$.
Let $h\coloneq v^{*}-v_{\sigma}$. By the definition of $v^{*}$ we have $0\leqslant h$. We claim in addition that $h=0$. To see this, suppose to the contrary that $h$ is nonzero. In this case, by strong irreducibility, for each nonzero $\mu$ in the positive cone of $b\mathsf{X}^{\prime}$ we can find an $m\in\mathbbm{N}$ such that $\left\langle{\mu,P_{\sigma}^{m}\,h}\right\rangle>0$. Because $\delta_{x}$ is a nonzero element of the positive cone of $b\mathsf{X}^{\prime}$ (Lemma 2.1), we can set $\mu=\delta_{x}$ to obtain an $m\in\mathbbm{N}$ with $(P_{\sigma}^{m}\,h)(x)>0$. This contradicts (6), so $h=0$ holds. In other words, $v_{\sigma}=v^{*}$. This proves (E3). ∎
@@ 4.1. Reachable States
As a preliminary step, we show that $h\coloneq v^{*}-v_{\sigma}$ is continuous under the stated assumptions. Since $\sigma$ is continuous, our conditions on $(r,\Gamma,\beta,P)$ imply that the mappings $x\mapsto\int v(x^{\prime})P(x,\sigma(x),dx^{\prime})$ and $x\mapsto r(x,\sigma(x))$ are continuous on $\mathsf{X}$ whenever $v\in cb\mathsf{X}$. This implies that $T_{\sigma}$ is invariant on $cb\mathsf{X}$. Moreover, $cb\mathsf{X}$ is a closed subset of the complete metric space $b\mathsf{X}$ under the supremum norm (since uniform limits of continuous functions are continuous). In addition, given $v,w\in b\mathsf{X}$, we have
@@
Since $\beta<1$, the contraction mapping theorem implies that $T_{\sigma}^{n}w\to v_{\sigma}$ for every $w\in b\mathsf{X}$. If we now fix $w\in cb\mathsf{X}$ and use the fact that $T_{\sigma}$ is invariant on this set, we obtain a sequence $(T^{n}w)_{n\in\mathbbm{N}}$ converging to $v_{\sigma}$ and entirely contained in $cb\mathsf{X}$. As $cb\mathsf{X}$ is closed in $b\mathsf{X}$, this implies that $v_{\sigma}$ is in $cb\mathsf{X}$. In particular, $v_{\sigma}$ is continuous. As $v^{*}$ is also continuous (see Proposition 2.2), we see that $h$ is continuous.
@@
But $v_{\sigma}(x)=v^{*}(x)$, so this inequality contradicts Lemma 2.3. The contradiction proves Lemma 4.1. ∎
@@ 4.2. Open Set Irreducibility
In view of Lemma 3.1, it suffices to show (E1) implies (E3). So fix $x\in\mathsf{X}$ and suppose that $v_{\sigma}(x)=v^{*}(x)$. For any $y\in\mathsf{X}$, open set irreducibility implies that $y$ is $P_{\sigma}$-reachable from $x$. Hence, by Lemma 4.1, we have $v_{\sigma}(y)=v^{*}(y)$. In particular, (E3) holds. ∎
@@ 4.3. $\pi$-Irreducibility
Now we state a result for the weakly irreducible case. In the statement, $\sigma$ is any feasible policy.
@@
5. Applications
In this section, we consider three applications. The first involves an optimal savings problem with stochastic returns on wealth, and our aim is to illustrate the significance of our main theoretical results. The second application provides an extension showing how results in the paper can be applied outside the MDP framework. The third application, found in Section 5.4, demonstrates that the irreducibility assumptions in Theorems 3.3, 4.2, and 4.4 cannot be removed without changing their conclusions.
5.1. Optimal Savings
Consider an agent who seeks to maximize lifetime utility by choosing optimal savings and consumption. The evolution of wealth is governed by the equation
@@ 5.1. Optimal Savings
We formulate this problem as an MDP. The state space is $\mathbbm{R}_{+}$ and the set of feasible actions at wealth level $w$ is $\Gamma(w)=\{c\in\mathbbm{R}_{+}:c\leqslant w\}$. A feasible policy in this setup is a Borel measurable function $\sigma$ from $\mathbbm{R}_{+}$ to itself satisfying $\sigma(w)\leqslant w$ for all $w\in\mathbbm{R}_{+}$. The reward function is $r(w,c)\coloneq u(c)$, where $u(c)$ is the utility derived from consumption and $u$ is continuous, differentiable, and strictly concave on $\mathbbm{R}_{+}$.
@@
Let $\Sigma$ be the set of feasible policies. Fix $\bar{w}\in\mathbbm{R}_{+}$ and an open set $B\subseteq\mathbbm{R}_{+}$ with $\pi(B)>0$. Let $\alpha=w-\sigma(w)$. By a change of variable, we obtain
@@
Given the open set irreducibility of the transition kernel at any feasible policy stated in Lemma 5.1, Theorem 4.2 implies that we can compute a (globally) optimal policy by maximizing $v_{\sigma}(\bar{w})$ at any fixed $\bar{w}\in\mathbbm{R}_{+}$. We now explore this result in a computational experiment, where the maximization of $v_{\sigma}(\bar{w})$ is based on a simple gradient ascent algorithm with a neural network to approximate the policy function $\sigma$.
Let $\theta\in\mathbbm{R}^{d}$ be the network parameters that characterize the policy network $\hat{\sigma}(\cdot;\theta)$ with a fixed paramter size $d\in\mathbbm{N}$. We want to find the policy parameterization that maximizes the expected discounted sum of rewards defined by $u(\hat{\sigma}(w;\theta))$ starting from the initial wealth level $\bar{w}$:
@@
$\max_{\hat{\sigma}\in\Sigma}v_{\hat{\sigma}}(\bar{w})\coloneq\max_{\hat{\sigma}\in\Sigma}\sum_{t=0}^{\infty}\beta^{t}\int u(\hat{\sigma}(w^{\prime};\theta))P^{t}_{\hat{\sigma}}(\bar{w},\mathop{}\!\mathrm{d}w^{\prime}).$
We can approximate this by Monte Carlo with large batch size $N$ and rollout steps $T$:
$\hat{v}_{\hat{\sigma}}(\bar{w})\coloneq\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\beta^{t}u(\hat{\sigma}(w_{i,t};\theta))\quad\text{with}\quad w_{i,0}=\bar{w}\text{ for }i=1,\ldots,N,$ (11)
Here $w_{i,t}$ is the wealth at time $t$ for the $i$-th sample, conditional on starting at $\bar{w}$ and following policy $\hat{\sigma}$. (In particular, we generate the rollout trajectories $(w_{i,t})_{t=0}^{T}$ for all $i=1,\ldots,N$ according to the wealth dynamics (9) with consumption at time $t$ following $\hat{\sigma}(w_{i,t};\theta)$.) The fact that $w_{i,0}=\bar{w}$ for all $i=1,\ldots,N$ means that we are optimizing lifetime value over policies while freezing the initial state $w_{0}=\bar{w}$.
During each episode, the model parameters are updated to reduce the value of the loss function
$L(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\beta^{t}u(\hat{\sigma}(w_{i,t};\theta))$ (12)
using gradient descent. It is clear that minimizing the loss is equivalent to maximizing the finite time approximation of the expected lifetime value (11).
The complete algorithm is detailed in Algorithm 1 in the Appendix. Our approach is similar to the REINFORCE algorithm (Williams, 1992), with a key distinction: since the utility function in our setup is differentiable, we can compute well-defined gradients without requiring a stochastic policy.
Several techniques could enhance the algorithm’s stability and convergence, including reinforcement learning baselines (Williams, 1992) and more advanced approaches applying actor-critic methods (Silver et al., 2014; Lillicrap et al., 2019). We deliberately avoid these extensions in this paper to maintain algorithmic clarity and simplicity, allowing us to establish a more direct connection to our theoretical results.
Figure 1. $v_{\hat{\sigma}}$ and $\hat{\sigma}$ with $\bar{w}=1$ vs OPI solutions
@@
Figure 2. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the irreducible optimal saving model at $\bar{w}=1$
To verify our theory, we benchmark the result from optimizing the policy at one point against Optimistic Policy Iteration (OPI), a variant of Value Function Iteration (VFI) that is known to converge globally to the optimal policy in this model-based setting (Sargent and Stachurski, 2025a). The OPI algorithm operates on a discretized state space and serves as our ground truth comparison. Specifically, we compute $v^{*}$ by applying OPI over a fine grid of wealth levels, which yields a global approximation of the optimal value function and corresponding optimal policy $\sigma^{*}$.
Regarding $v_{\hat{\sigma}}$, we compute this function in two steps. First, we freeze a single initial wealth level $\bar{w}$ and minimize the loss according to (12), as described above. Second, we take the resulting policy $\hat{\sigma}$ and, holding this policy fixed, calculate the function $v_{\hat{\sigma}}$ by computing the lifetime value of $\hat{\sigma}$ from alternative initial conditions. Finally, we compare $v_{\hat{\sigma}}$ with the globally optimal solution $v^{*}$. In the experiments, we use the CRRA utility function $u(c)=c^{1-\gamma}/(1-\gamma)$ with $\gamma=2$.
@@
Figure 2 shows the result of these computations when $\bar{w}=1$. The function $v_{\hat{\sigma}}$, shown in blue, closely matches the globally optimal value function $v^{*}$ computed via OPI (red dotted line). Similarly, in the second panel, the approximated policy $\hat{\sigma}$ (blue line) closely matches the optimal policy $\sigma^{*}$ (red dotted line). This convergence demonstrates that both the $\hat{\sigma}$-value function $v_{\hat{\sigma}}$ and the policy $\hat{\sigma}$ successfully recover their globally optimal counterparts $v^{*}$ and $\sigma^{*}$, respectively. This is consistent with the result in Theorem 4.2, which states that we can compute a globally optimal policy by solving $\max_{\sigma\in\Sigma}v_{\sigma}(\bar{w})$ at any fixed $\bar{w}\in\mathbbm{R}_{+}$.
@@
In subsequent experiments, we tested the robustness of this outcome to variation in the fixed initial condition $\bar{w}$. We found that, as predicted by theory, the resulting function $v_{\hat{\sigma}}$ again closely approximates $v^{*}$, and the resulting policy $\hat{\sigma}$ again closely approximates $\sigma^{*}$. See Figure 7 in Appendix B for a visualization of one experiment.
In Figure 2, the $y$-axis is the finite time approximation of the infinite horizon lifetime value function $v_{\hat{\sigma}}$ defined in (11) and the $x$-axis is the number of training episodes, which records the number of times the policy $\hat{\sigma}$ is updated according to the $\hat{v}_{\hat{\sigma}}(\bar{w})$ computed from the $N$ Monte Carlo samples. The figure shows consistent improvement in approximated lifetime value from following the policy $\hat{\sigma}$ that improves over the training episodes. The sequence of approximated lifetime value $\hat{v}_{\hat{\sigma}}(\bar{w})$ exhibit a sharp increase in the first half of the training episodes, followed by a stable plateau.  [^2] The sequence of approximated lifetime value converges to a higher value than $v_{\hat{\sigma}}(1)$. This is because the $\hat{v}_{\hat{\sigma}}(\bar{w})$ is a finite time approximation (with rollout $T$) of the infinite horizon lifetime value function $v_{\hat{\sigma}}$. The same phenomenon occurs in Figure 7 and subsequent experiments in the reducible case.
5.2. A Reducible MDP